OpenStudy (anonymous):

17. Simplify the following: 4x – 2 – 3x 4x – (2 – 3x)

OpenStudy (anonymous):

4x - 2 - 3x = x - 2 4x - (2 - 3x) = 4x - 2 + 3x = 7x - 2